# Magnetoconvection in a spherical shell: Equatorial symmetry during the transition from the weak- to the strong-field regime

### 09 Mar 2026 v1

@@ 1 Introduction
Dynamo simulations exhibit a wide range of solutions. Some are essentially kinematic, with the magnetic field generated by—yet only weakly influencing—the flow dynamics. Other states emerge only once the magnetic field becomes sufficiently strong. In Earth‑like simulations, three principal branches are now recognised, each associated with distinct magnetic‑field morphologies and strengths (Dormy et al., 2018; Dormy, 2025). These states are commonly characterised by the dominant balance between physical forces, or more precisely, in incompressible flows, by the balance between the solenoidal components of those forces (Hughes and Cattaneo, 2019; Teed and Dormy, 2023), thereby excluding pressure, which acts as a Lagrange multiplier. Although the force balance in Earth’s core is well understood, it cannot be reproduced exactly in simulations. Because force balances underpin much of the scaling and structure of dynamo behaviour (Christensen, 2010), accurately identifying subtle shifts in these balances—and their consequences for flow and field morphology—is essential for reliable extrapolation from simulations to geophysical conditions.
@@
Nonlinear geodynamo simulations are often complemented by linear convection studies, which are computationally inexpensive and therefore provide a broader view of the families of modes that can arise. In differentially heated spheres and spherical shells, geostrophic modes onset when thermal or compositional forcing overcomes the stabilising effects of rotation and gravity; viscosity also plays a crucial role in mediating the Taylor–Proudman constraint (Roberts, 1968; Busse, 1970). The relative importance of viscous and Coriolis forces is quantified by the Ekman number, ${\rm Ek}=\nu/(\Omega L^{2})$, where $\nu$ is the kinematic viscosity, $\Omega$ the rotation rate, and $L$ a characteristic length scale. Near convective onset, viscosity is the primary mediator of rotational constraints but becomes significant only at small scales; consequently, the large‑scale flow is predominantly $z$-independent and takes the form of precessing columnar Rossby waves with heights of the order of the shell and widths scaling as ${\rm Ek}^{-1/3}$, consistent with the leading‑order VAC (viscous–Archimedean–Coriolis) balance. A single discrete mode typically dominates when its characteristic scale is large, suppressing the linear growth of competing modes. The precessing columns bend radially inward around the inner core, enhancing radial heat transport (Busse, 2002). Their columnar rotation, together with a thermally driven axial flow, can drive an $\alpha^{2}$-dynamo provided the shearing rate exceeds the rate of magnetic diffusion, usually quantified by the magnetic Reynolds number. In the VAC regime, magnetic diffusivity does not enter the leading‑order force balance and its variation does not significantly alter the flow structure. Thus, one path to achieve a weak dynamo is to reduce magnetic diffusivity (i.e., increase the magnetic Prandtl number), though this can substantially modify the magnetic‑field properties (Teed and Dormy, 2025a).
@@
In the presence of a strong magnetic field, magnetic tension resists fluid motion that bends field lines, generally stabilising the fluid against convection. However, in rapidly rotating systems, a weak magnetic field modifies the Taylor-Proudman theorem to allow fully three-dimensional convective (magnetostrophic) modes to onset with a significantly lower critical Rayleigh number, reaching a minimum when the ratio of the Lorentz and Coriolis forces, the Elsasser number, is of order one (Chandrasekhar, 1961; Fearn, 1979). Geostrophic and magnetostrophic modes can have similar Rayleigh numbers at Elsasser numbers only just below one; when these modes coexist, the wider array of potential convective modes is thought to allow strong-field dynamos, which are essentially chaotic, with continuous turnover in the prevalence of several modes.
@@
Sakuraba (2002) shows that polar convective modes, that is, modes which peak in amplitude inside the tangent cylinder, can have critical Rayleigh numbers similar to the geostrophic and magnetostrophic modes at order unity Elsasser numbers. Polar convective modes are distinct in that they could onset at critical Rayleigh numbers close to the values found in linear theory when convection is present, but significant only in the bulk (Gilman, 1975, 1977). Sreenivasan and Jones (2006) use an $f$-plane model to study these modes and show that they onset as axial plumes with oppositely signed vorticity at their upper and lower ends. Unlike (magneto)geostrophic modes, the dominant component of the flow is axial, directly supporting the poloidal magnetic field.
@@
In a spherical shell, polar convective modes can onset with either equatorially symmetric or antisymmetric eigenfunctions with no strong preference for either (Appendix B). Equatorial symmetry is the preferred onset configuration for modes in the bulk, due to the Taylor-Proudman constraint (Busse, 1970; Sakuraba, 2002), and this symmetry is preserved by the nonlinear terms of the momentum and induction equations. The breaking of equatorial symmetry can thus be caused by the onset of polar convective modes, and the change in symmetry will significantly alter the structure of the flow. In particular, the breaking of equatorial symmetry can allow for the formation of large scale meridional circulation currents that can strengthen the poloidal magnetic field and support its reversals.
@@
Astrophysical dynamos lie further into the diffusion-free asymptotic realm than can be reached in numerical simulation. Many studies push towards low $\rm Ek$}, $\rm Pm$}, seeking to derive scaling laws independent of $\nu$, $\eta$, and $\kappa$, however, a refinement of this method is to push towards this asymptotic regime in an ordered way, along a distinguished limit that preserves a desired property of the simulation, such as the force balance, magnetic field morphology, or reversals. To capture these properties, at moderate Ekman numbers, it is often necessary to use larger than physical values of the magnetic or thermal Prandtl numbers (i.e., weaker than physical thermal or magnetic diffusion, e.g., Dormy, 2016; Jones and Tsang, 2025). However, at these larger values of the magnetic Prandtl number, $\rm Pm$}, both weak-field (columnar) and strong-field solutions can co-exist at the same input parameters. This bistability between dynamo branches means that the resultant state depends on the initial state of the flow (Dormy, 2016; Teed and Dormy, 2025a) (Bistability has also been shown to occur between the strong-field and inertia-dominated branches; Simitev and Busse, 2009; Menu et al., 2020).
@@
One possible way to circumvent this problem of bistability is to modify the boundary conditions on the magnetic field so that it is either strengthened or weakened. The resulting simulation corresponds to magnetoconvection rather than a dynamo, but the properties of the flow are generally similar, provided that the imposed boundary conditions are consistent with those inferred from equivalent dynamo simulations (§3.2). We find that bistability occurs only for moderately strong imposed fields (4). Another advantage of magnetoconvection is that it allows for the direct determination of the influence of the magnetic field on flow, for example, by promoting convective modes with larger length scales when magnetic stresses replace viscosity as the primary mechanism that breaks the Taylor–Proudman constraint.
@@
In this work, we use magnetoconvection simulations to study the weak- to strong-field transition. We also focus on the phenomenon of symmetry-breaking, which coincides with this transition. In §2.3, we outline the governing equations, our choice of parameterisation, and define several important output quantities, including a symmetry parameter, $s_{U/B}$, which measures the symmetry of the flow and magnetic field respectively. We then review suites of hydrodynamic and dynamo simulations from (Teed and Dormy, 2025b, a) in §3, to understand how these output quantities vary in those simulations and provide context for the results of §4, in which we present magnetoconvection simulations at three values of $\rm Pm$}, chosen to portray a wide array of possible phase spaces, with respect to the onset of the weak- and strong-field branches. Finally, we summarise in §5 and, based on our results, outline a possible mechanism by which equatorial symmetry is broken.
@@ 2.1 Mathematical and numerical setup
With $\rm Pr$}, $\rm Ek$}, and $\chi$ fixed, we will present results from three types of simulation: The first are hydrodynamic simulations, the equation for which can be derived from (1a)-(1c) with $\boldsymbol{B}\equiv\boldsymbol{0}$ and ${\rm Pm}=1$, where the latter criterion changes the effective nondimensionalisation, and the second are dynamo simulations for which electrically-insulating boundary conditions are used (Appendix A); these two types of simulation correspond to the same dataset as presented by Teed and Dormy (2025a, b). These are presented in §3 as motivation for and comparison with the main set of simulations: magnetoconvection simulations, in which the radial component of the magnetic field is fixed at the outer boundary (CMB), and the inner boundary (ICB) is kept with insulating conditions. Details are provided in Appendix A. The range of appropriate values for the dipolar magnetic field strength at the outer boundary, $p_{10}$, are chosen by comparison with the dynamo simulations and are detailed at the end of §3. In our dynamo and magnetoconvection simulations, magnetic Prandtl numbers in the range $[1,12]$ are used, which, given the other input parameters, covers a region of parameters space that allows for different bifurcations and solution branches in dynamo solutions Teed and Dormy (2025a). In particular, large Pm values in this range admit separate weak- and strong-field dipolar branches with bistability between them and such solutions correspond to those envisaged for the distinguished limit of Dormy (2016).
@@ 2.2 Output quantities
$\frac{|\left(\boldsymbol{j}\times\boldsymbol{B}\right)|}{2\rho\left|\boldsymbol{\Omega}\times\boldsymbol{u}\right|}\sim\Lambda=\frac{\left<B\right>^{2}}{2\rho\mu\eta\Omega},$ (5)
@@
$\Lambda^{\prime}=\frac{\left<B\right>^{2}}{2\Omega\mu\rho\left<U\right>\ell_{\eta}}=\Lambda\frac{L}{{\rm Rm}\ell_{\eta}},$ (6)
@@
$\ell_{\eta}^{2}=\frac{\left<\boldsymbol{B}\right>^{2}}{\left<\boldsymbol{j}\right>^{2}}$ (7)
@@
In our magnetoconvection work, since the magnetic field is partly controlled at the boundary and may be artificially strong or weak (as required), we introduce a modified Elsasser measure,
$\Lambda^{\prime}_{\neg as}=\frac{\left<B\right>_{m\neq 0}^{2}}{2\Omega\mu\rho\left<U\right>\ell_{\eta}}\leq\Lambda^{\prime},$ (8)
in which the axisymmetric component of the magnetic energy is eliminated. We find that $\Lambda^{\prime}_{\neg as}$ provides a useful method for measuring the influence of the small-scale magnetic field in magnetoconvection simulations, especially when used in concert with $\Lambda^{\prime}$ and the known input magnetic field strength.
A key change in the flow at weak nonlinearity as thermal forcing is increased and convection becomes more vigorous is the breaking of the Taylor-Proudman constraint, that is, the trend away from a largely $z$-invariant flow. Soderlund et al. (2012) suggest that the anisotropy of such a flow can be measured by its columnarity, the relative size and $z$-independence of the $z$-component of the vorticity. We calculate this as
@@
$\boldsymbol{u}_{s}=\frac{\boldsymbol{u}\left(\theta\right)+\boldsymbol{u}^{r}\left(\pi-\theta\right)}{2}\qquad\boldsymbol{u}_{as}=\frac{\boldsymbol{u}\left(\theta\right)-\boldsymbol{u}^{r}\left(\pi-\theta\right)}{2},$ (10a,b)
@@
$\text{KE}_{\perp}$ $=\frac{1}{2}\left<\left|\boldsymbol{u}\right|\right>^{2}-\text{KE}_{zonal}-\text{KE}_{\parallel}\,$, (11c)
@@ 2.3 Minimal field configuration
$\text{ME}^{*}=\rho\eta^{2}D\frac{{\rm Pm}}{{\rm Ek}}\int_{V}\frac{B^{2}}{2}\penalty 10000\ {\rm d}V,$ (15)
@@ 3.1 Weakly-nonlinear hydrodynamic convection
where $U_{\parallel}/U_{\perp}\approx\sqrt{0.2}$ can be determined from the energy balance (Figure 1(c)), and we can take $\ell_{\parallel}=2\sqrt{r_{o}^{2}-s_{c}^{2}}$, the height of the sphere at the critical radius, $s_{c}\approx 0.42r_{o}$, and $\ell_{\perp}=2s_{c}\pi/m_{c}$, where $m_{c}=7$ is the mode of onset. Using these, we can obtain an estimate of $C_{\omega z}\approx 0.955$, which is slightly larger than the value calculated directly. As the Rayleigh number increases, the flow becomes unstable further away from the critical radius and the convective columns increase their radial extent, slightly increasing $\ell_{\perp}$, and therefore slightly decreasing $C_{\omega z}$. The onset modes are equatorially symmetric (Busse, 1970), and hence the flow remains equatorially symmetric since the nonlinear terms preserve this symmetry (Figure 1(b)); antisymmetric modes must onset separately.
@@
Increasing the Rayleigh number above ${\rm Ra}^{\prime}=2$ eventually leads to an instability of the axial columns which results in a steady decline in $C_{\omega z}$; this also allows other modes to grow and eventually dominate over the primary onset mode. The columnarity, Figure 1(a), can also be compared with Figure 3a of Soderlund et al. (2012), who use $\chi=0.4$, i.e. a slightly larger inner core radius, and ${\rm Ek}\approx 1.7\times 10^{-4}$ [^2] Soderlund et al. (2012) use ${\rm Ek}=10^{-4}$, but define their Ekman number as ${\rm Ek}=\nu/2\Omega D^{2}$, where $D$ is the distance between the outer and inner boundaries and also use $\chi=0.4$; one way to convert is to fix $r_{o}$, the outer radius, which gives the scale factor $2\times(1-0.4)^{2}/(1-0.35)^{2}\approx 1.7$.. This leads to a more rapid decrease in $C_{\omega z}$ with $\rm Ra$$^{\prime}$ since the preferred azimuthal wavenumber of convection, $m_{c}$, grows due to the larger inner core (e.g. Seeley et al., 2025), and hence convective columns become unstable to centrifugal instabilities at lower $\rm Ra$$^{\prime}$.
@@
At ${\rm Ra}^{\prime}\approx 10$ the symmetry, $s_{u}$, suddenly decreases due to the convective onset of equatorially antisymmetric modes. Gilman (1977) describes a similar situation for the solar convection envelope; since the eigenfunction of the equatorially symmetric onset mode, and hence the resultant convection columns, are localised to the region outside the tangent cylinder, the polar modes, whose eigenfunctions peak inside the tangent cylinder, onset at approximately the same Rayleigh number as expected from linear theory; meanwhile, the non-dominant equatorial modes are strongly suppressed by the $m_{c}=7$ mode and onset well beyond the linear value. Gastine and Aurnou (2023) have also raised the importance of the polar mode, discussing it in the context of significantly increasing $\rm Nu$}, the heat flux; this occurs since convection cells can develop in the tangent cylinder whose radial flow is less affected by the Coriolis force. We shall see later that this increased radial flow also affects the generation of the poloidal field. The lack of a jump in $C_{\omega z}$ at ${\rm Ra}^{\prime}=10$, which, unlike $s_{u}$, is calculated only over the region outside the tangent cylinder, supports the idea that the antisymmetric instability onsets inside the tangent cylinder.

### 17 Mar 2026 v2

@@ 1 Introduction
Dynamos, or self-sustaining magnetic fields, are generated by the convective motions of electrically-conducting fluids in astrophysical bodies and ubiquitous in the Universe; simulations of these systems exhibit a wide range of solutions. Some are essentially kinematic, with the magnetic field generated by—yet only weakly influencing—the flow dynamics. Other states emerge only once the magnetic field becomes sufficiently strong. In simulations modelling Earth’s core, three principal branches are now recognised, each associated with distinct magnetic‑field morphologies and strengths (Dormy et al., 2018; Dormy, 2025). These states are commonly characterised by the dominant balance between physical forces, or more precisely, in incompressible flows, by the balance between the solenoidal components of those forces, thereby excluding pressure, which acts as a Lagrange multiplier (Hughes and Cattaneo, 2019; Teed and Dormy, 2023). The expected force balance in Earth’s core cannot be reproduced exactly in simulations due to a lack of adequate computing power. Because force balances underpin much of the scaling and structure of dynamo behaviour (Christensen, 2010), accurately identifying subtle shifts in these balances—and their consequences for flow and field morphology—is essential for reliable extrapolation from simulations to geophysical conditions.
@@
Nonlinear geodynamo simulations are often complemented by linear convection studies, which are computationally inexpensive and therefore provide a broader view of the families of modes that can arise. In differentially heated spheres and spherical shells, geostrophic modes onset when thermal or compositional forcing overcomes the stabilising effects of rotation and gravity; viscosity also plays a crucial role in mediating the Taylor–Proudman constraint, which demands invariance of flows along the rotation axis (Roberts, 1968; Busse, 1970). The relative importance of viscous and Coriolis forces is quantified by the Ekman number, ${\rm Ek}=\nu/(\Omega L^{2})$, where $\nu$ is the kinematic viscosity, $\Omega$ the rotation rate, and $L$ a characteristic length scale. Near convective onset, viscosity is the primary mediator of rotational constraints but becomes significant only at small scales; consequently, the large‑scale flow is predominantly axially-independent and takes the form of precessing columnar Rossby waves with heights of the order of the shell and widths scaling as ${\rm Ek}^{-1/3}$, consistent with the leading‑order VAC (viscous–Archimedean–Coriolis) balance. A single discrete mode typically dominates when its characteristic scale is large, suppressing the linear growth of competing modes. The precessing columns bend radially inward around the inner core, enhancing radial heat transport (Busse, 2002). Their columnar rotation, together with a thermally driven axial flow, can drive an $\alpha^{2}$-dynamo provided the shearing rate exceeds the rate of magnetic diffusion, usually quantified by the magnetic Reynolds number (e.g. Proctor and Gilbert, 1994). In the VAC regime, the Lorentz force is absent from the leading-order force balance so variation of the magnetic field does not significantly alter the flow structure. Thus, one path to achieve a weak dynamo is to reduce magnetic diffusivity $\eta$ (i.e. increase the magnetic Prandtl number, ${\rm Pm}=\nu/\eta$), though this can substantially modify the magnetic‑field properties (Teed and Dormy, 2025a).
@@
In the presence of a strong magnetic field, magnetic tension resists fluid motion that bends field lines, generally stabilising the fluid against convection; that is, the critical Rayleigh number, which measures the (relative) strength of thermo-compositional forcing required for convective onset, is large. However, in rapidly rotating systems, a magnetic field for which the ratio of the Lorentz and Coriolis forces, the Elsasser number, is of order one, modifies the Taylor-Proudman constraint to allow fully three-dimensional convective (magnetostrophic) modes to onset at significantly small critical Rayleigh numbers (Chandrasekhar, 1961; Fearn, 1979). Geostrophic and magnetostrophic modes can have similar critical Rayleigh numbers at Elsasser numbers only just below one (Horn and Aurnou, 2022); when these modes coexist, the wider array of potential convective modes, with sufficient nonlinearity, can lead to chaotic flows that may promote strong-field dynamos.
@@
The tangent cylinder is the cylindrical surface aligned with the rotation axis that circumscribes the inner core. Dynamics inside the tangent cylinder are often disconnected from dynamics outside the tangent cylinder due to the effect of the Taylor-Proudman constraint. Sakuraba Sakuraba (2002) shows that polar convective modes, that is, modes which peak in amplitude inside the tangent cylinder, can have critical Rayleigh numbers similar to the geostrophic and magnetostrophic modes at order unity Elsasser numbers. Polar convective modes may also onset in nonlinear simulations at critical Rayleigh numbers close to values found in linear theory, when convection is significant only outside the tangent cylinder (Gilman, 1975, 1977). Sreenivasan & Jones Sreenivasan and Jones (2006) use an $f$-plane model to study these modes and show that they onset as axial plumes with oppositely signed vorticity at their upper and lower ends. Unlike equatorial modes, the dominant flow component of a polar convective mode is axial, directly supporting the poloidal magnetic field.
@@
In a spherical shell, polar convective modes can onset with either equatorially symmetric or antisymmetric eigenfunctions with no strong preference for either (Appendix B). Equatorial symmetry is the preferred onset configuration for modes outside the tangent cylinder, due to the Taylor-Proudman constraint (Busse, 1970; Sakuraba, 2002), and this symmetry is preserved by the nonlinear terms of the momentum and induction equations. The breaking of equatorial symmetry can thus only be caused by the separate onset of antisymmetric modes, such as polar convective modes, and the change in symmetry will significantly alter the structure of the flow Christensen et al. (1999). In particular, the breaking of equatorial symmetry can allow for the formation of large scale meridional circulation currents that can strengthen the poloidal magnetic field and support its reversals (Sarson and Jones, 1999; Wicht and Olson, 2004).
@@
Astrophysical dynamos lie further into the diffusion-free asymptotic realm than can be reached in numerical simulation. Many studies push towards low $\rm Ek$}, $\rm Pm$}, seeking to derive scaling laws independent of $\nu$, $\eta$, and $\kappa$; however, a refinement of this method is to push towards this asymptotic regime in an ordered way, along a distinguished limit (Dormy, 2016) that preserves a desired property of the simulation, such as the force balance, magnetic field morphology, or reversals. To capture these properties, at moderate Ekman numbers, it is often necessary to use larger than physical values of the magnetic or thermal Prandtl numbers (i.e., weaker than physical thermal or magnetic diffusion, e.g., Dormy, 2016; Jones and Tsang, 2025). However, at these larger values of $\rm Pm$}, both weak-field (columnar) and strong-field dipolar solutions can co-exist at the same input parameters. This bistability between dynamo branches means that the resultant state depends on the initial state of the simulation with the strong-field regime emerging due to the triggering of non-geostrophic modes (Dormy, 2016; Teed and Dormy, 2025a). Separately, bistability has also been shown to occur between the weak-field dipolar and inertia-dominated multipolar branches (Simitev and Busse, 2009), although it is unclear whether this type of bistability is possible between the strong-field dipolar and multipolar branches (Menu et al., 2020).
@@
One possible way to circumvent the issue of bistability is to modify the boundary conditions on the magnetic field so that it is either strengthened or weakened (Teed et al., 2015; Mason et al., 2022). The resulting simulation corresponds to magnetoconvection rather than a dynamo, but the properties of the flow are generally similar, provided that the imposed boundary conditions are consistent with those inferred from equivalent dynamo simulations (§3.2). We find that bistability occurs only for moderately strong imposed fields (§4). Another advantage of magnetoconvection is that it allows for the direct determination of the influence of the magnetic field on flow, for example, by promoting convective modes with larger length scales when magnetic stresses replace viscosity as the primary mechanism that breaks the Taylor–Proudman constraint.
@@
In this work, we use magnetoconvection simulations to study the weak- to strong-field transition with a particular focus on the phenomenon of symmetry-breaking, which coincides with this transition. In §2.3, we outline the governing equations, our choice of parameterisation, and define several important output quantities, including a symmetry parameter, which measures the equatorial symmetry of the flow and magnetic field respectively. We then review suites of hydrodynamic and dynamo simulations from (Teed and Dormy, 2025b, a) in §3, to understand how these output quantities vary in those simulations and provide context for our magnetoconvection results. In §4, we present magnetoconvection simulations at three values of $\rm Pm$}, chosen to portray a wide array of possible phase spaces, with respect to the onset of the weak- and strong-field branches. Finally, we summarise in §5 and, based on our results, outline a possible mechanism by which equatorial symmetry is broken.
@@ 2.1 Mathematical and numerical setup
With $\rm Pr$}, $\rm Ek$}, and $\chi$ fixed, we will present results from three types of simulation: The first are hydrodynamic simulations, the equation for which can be derived from (1a)-(1c) with $\boldsymbol{B}\equiv\boldsymbol{0}$ and ${\rm Pm}=1$, where the latter criterion changes the effective nondimensionalisation, and the second are dynamo simulations for which electrically-insulating boundary conditions are used (Appendix A); these two types of simulation correspond to the same dataset presented by Teed and Dormy (2025a, b). These are presented in §3 as motivation for and comparison with the main set of simulations: magnetoconvection simulations, in which the radial component of the magnetic field is fixed at the outer boundary (CMB), and the inner boundary (ICB) is kept with insulating conditions. Details are provided in Appendix A. The range of appropriate values for the dipolar magnetic field strength at the outer boundary, $p_{10}$, are chosen by comparison with the dynamo simulations and are detailed at the end of §3. In our dynamo and magnetoconvection simulations, magnetic Prandtl numbers in the range $[1,12]$ are used, which, given the other input parameters, covers a region of parameters space that allows for different bifurcations and solution branches in dynamo solutions Teed and Dormy (2025a). In particular, large Pm values in this range admit separate weak- and strong-field dipolar branches with bistability between them, and such solutions correspond to those envisaged for Dormy’s distinguished limit Dormy (2016).
@@ 2.2 Output quantities
$\frac{|\left(\boldsymbol{\nabla}\times\boldsymbol{B}\right)\times\boldsymbol{B}|}{2\rho_{0}\mu_{0}\left|\boldsymbol{\Omega}\times\boldsymbol{u}\right|}\sim\Lambda=\frac{\left<B\right>^{2}}{2\rho_{0}\mu_{0}\eta\Omega},$ (5)
@@
$\Lambda^{\prime}=\frac{\left<B\right>^{2}}{2\Omega\mu_{0}\rho_{0}\left<U\right>\ell_{\eta}}=\Lambda\frac{L}{{\rm Rm}\ell_{\eta}},$ (6)
@@
$\ell_{\eta}^{2}=\frac{\left<\boldsymbol{B}\right>^{2}}{\left<\boldsymbol{\boldsymbol{\nabla}\times\boldsymbol{B}}\right>^{2}}$ (7)
@@
Since forces can vary across length scales, such as viscosity, which predominantly enters at small scales, so can the force balances. We make use of force spectra plots for which each force is plotted as a function of the spherical degree $l$ and averaged over a long time-scale. Following the discussion above, we use the curl of forces, which is numerically more straightforward to calculate than the solenoidal component, but still removes gradient terms. Forces are then divided by $\sqrt{l\left(l+1\right)}$ to remove the imbalance caused by the angular derivative - see Teed and Dormy (2023).
A key change in the flow at weak nonlinearity as thermal forcing is increased and convection becomes more vigorous is the breaking of the Taylor-Proudman constraint, that is, the trend away from a largely $z$-invariant flow. Soderlund et al.Soderlund et al. (2012) suggest that the anisotropy of such a flow can be measured by its columnarity, the relative size and $z$-independence of the $z$-component of the vorticity. We calculate this as
@@
$\boldsymbol{u}_{s}=\frac{\boldsymbol{u}\left(\theta\right)+\boldsymbol{u}^{r}\left(\pi-\theta\right)}{2}\qquad\boldsymbol{u}_{as}=\frac{\boldsymbol{u}\left(\theta\right)-\boldsymbol{u}^{r}\left(\pi-\theta\right)}{2},$ (9a,b)
@@
$\text{KE}_{\perp}$ $=\frac{1}{2}\left<\left|\boldsymbol{u}\right|\right>^{2}-\text{KE}_{\text{zonal}}-\text{KE}_{\parallel}\,$, (10c)
@@ 2.3 Minimal field configuration
$\text{ME}^{*}=\rho_{0}\eta^{2}D\frac{{\rm Pm}}{{\rm Ek}}\int_{V}\frac{B^{2}}{2}\penalty 10000\ {\rm d}V,$ (14)
@@ 3.1 Weakly-nonlinear hydrodynamic convection
where $U_{\parallel}/U_{\perp}\approx\sqrt{0.2}$ can be determined from the energy balance (Figure 1(c)), and we can take $\ell_{\parallel}=2\sqrt{r_{o}^{2}-s_{c}^{2}}$, the height of the sphere at the critical radius, $s_{c}\approx 0.42r_{o}$, and $\ell_{\perp}=2s_{c}\pi/m_{c}$, where $m_{c}=7$ is the mode of onset. Using these, we can obtain an estimate of $C_{\omega z}\approx 0.955$, which is slightly larger than the value calculated directly. As the Rayleigh number increases, the flow becomes unstable further away from the critical radius and the convective columns increase their radial extent, slightly increasing $\ell_{\perp}$, and therefore slightly decreasing $C_{\omega z}$. The preferred onset modes are equatorially symmetric (Busse, 1970), and hence the flow remains equatorially symmetric since the nonlinear terms preserve this symmetry (Figure 1(b)); antisymmetric modes must onset separately.
@@
Increasing the Rayleigh number above ${\rm Ra}^{\prime}=2$ eventually leads to an instability of the axial columns which results in a steady decline in $C_{\omega z}$; this also allows other modes to grow and eventually dominate over the primary onset mode. The columnarity, Figure 1(a), can also be compared with Figure 3a of Soderlund et al. (2012), who use $\chi=0.4$, i.e. a slightly larger inner core radius, and ${\rm Ek}\approx 1.7\times 10^{-4}$ [^2] Soderlund et al. (2012) use ${\rm Ek}=10^{-4}$, but define their Ekman number as ${\rm Ek}=\nu/2\Omega D^{2}$, where $D$ is the distance between the outer and inner boundaries and also use $\chi=0.4$; one way to convert is to fix $r_{o}$, the outer radius, which gives the scale factor $2\times(1-0.4)^{2}/(1-0.35)^{2}\approx 1.7$.. This leads to a more rapid decrease in $C_{\omega z}$ with $\rm Ra$$^{\prime}$ since the preferred azimuthal wavenumber of convection, $m_{c}$, grows due to the larger inner core (e.g. Seeley et al., 2025), and hence the convective columns become unstable to centrifugal instabilities at lower $\rm Ra$$^{\prime}$.
@@
At ${\rm Ra}^{\prime}\approx 10$ the symmetry, $s_{u}$, suddenly decreases due to the convective onset of equatorially antisymmetric modes. Gilman Gilman (1977) describes a similar situation for the solar convection envelope; since the eigenfunction of the equatorially symmetric onset mode, and hence the resultant convection columns, are localised to the region outside the tangent cylinder, the polar modes, whose eigenfunctions peak inside the tangent cylinder, onset at approximately the same Rayleigh number as expected from linear theory; meanwhile, the non-dominant equatorial modes are strongly suppressed by the $m_{c}=7$ mode and onset well beyond the linear value. Gastine & Aurnou Gastine and Aurnou (2023) have also raised the importance of the polar mode, discussing it in the context of significantly increasing $\rm Nu$}, the heat flux; this occurs since convection cells can develop in the tangent cylinder whose radial flow is less affected by the Coriolis force. We shall see later that this increased radial flow also affects the generation of the poloidal field. The lack of a jump in $C_{\omega z}$ at ${\rm Ra}^{\prime}=10$, which, unlike $s_{u}$, is calculated only over the region outside the tangent cylinder, supports the idea that the antisymmetric instability onsets inside the tangent cylinder.